Question 2. (10 points) Suppose the demand and supply curves for units of university credits are given by the following equations: Q D = 5000 − P Q S = 3P − 200 where QD is the quantity of credits demanded, QS is the quantity supplied, and P is the price charged for each unit in dollars.

(a) (3 points) What is the free-market equilibrium Price and Quantity. 3

(b) (3 points) Suppose that the government wants to make education more accessible and therefore passes a regulation that says no university can charge more than $800 per credit. What is the new equilibrium Price and Quantity?

(c) (4 points) Calculate the shortage or surplus generated by this government policy. Show it on graph. Include your answers from part (a) and (b) on the graph.
Answer #1

a) Set QD=QS

5000-P = 3P-200

5000+200 = 3P+P

P = 5200/4 = 1300

Q = 5000-1300 = 3700

b) When the government sets the maximum price at P=800

QD = 5000-800 = 4200

QS = 3*800-200 = 2200

New equilibrium price = 800

New equilibrium quantity = 2200

c) There would be a shortage of 4200-2200 = 2000 units

1300 800 Price ceiling 2200 3700 4200